mhchen:

Solve x^2 - 3x + 1 = 0 algebraically

mhchen:

I used formula: \[ax^{2}+bx+c = a(x-\frac{-b-\sqrt{b^{2}-4ac}}{2a})(x-\frac{-b+\sqrt{b^2{}-4ac}}{2a}) = 0\] Which is pretty cheap and lame. Is there a better/simpler way to solve it?

dude:

What do you mean "better and simpler" If you're looking for another way to solve it just complete the square, if you're looking for a fancy way, then I don't know.

mhchen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude What do you mean "better and simpler" \(\color{#0cbb34}{\text{End of Quote}}\) Like usually I can rewrite it in (x-b)(x-d) = 0 by looking at it but I couldn't for this one. I will try completing the square. \[x^{2}-3x+(\frac{3}{2})^{2}=-1+(\frac{3}{2})^{2}\] \[(x-\frac{3}{2})^2 = \frac{4}{4}+\frac{9}{4}\] \[x-\frac{3}{2} = \pm \sqrt{\frac{13}{4}}\] \[x = \frac{3}{2}+\sqrt{\frac{13}{4}},x=\frac{3}{2}-\sqrt{\frac{13}{4}}\] lol it works